# Gravitational Waves Emission in Quadratic Gravity: longitudinal modes, angular momentum emission, and positivity of the radiated power

### v1

@@ IV Rotating ellipsoid with free precession
To analyze the impact of massive modes on rotational dynamics, we will focus our study specifically on the tensor mode within the oscillatory regime. In this case, by integrating Eqs. (199) and (200), we get the curves that are shown in Figs. 3 and 4. Figure 3 clearly shows that the frequency decreases with time, similar to what is predicted by GR. However, it also shows that the inclusion of a massive field moderates this frequency reduction, causing it to decrease more slowly as the mass $m_{\Psi}$ decreases. This happens because the reduction in $m_{\Psi}$ increases the destructive interference between the fields $h_{\mu\nu}$ and $\Psi_{\mu\nu}$, leading to a slower loss of energy by the system. A similar effect was observed in [40] for point-like binary mass systems. The greater effectiveness of destructive interference with decreasing mass can be understood by recalling that $m_{\Psi}\propto\frac{1}{\alpha}$, where $\alpha$ is the coupling constant associated with the $C^{2}$ term in the action. Therefore, as $m_{\Psi}$ decreases, $\alpha$ increases, amplifying the impact of the massive spin-$2$ field.
@@ V Final comments
In this paper, we analyzed the emission of gravitational waves within the framework of quadratic gravity, including the contribution of longitudinal and transverse massive modes, extending the previous analysis presented in Ref. [40]. Firstly, we derived the wave equations for the massless ($\tilde{h}_{\alpha\beta}$) and massive ($\Phi,\Psi_{\alpha\beta}$) perturbations, followed by obtaining the associated conserved currents. Subsequently, we calculated the energy and angular momentum radiated through gravitational waves. Focusing on the quadrupole approximation, we examined two particular cases: a binary point-mass system in circular orbit, and a rotating ellipsoid with free precession.
@@ Acknowledgements
MFSA thanks FAPES (Brazil) for its support. The authors RRC and LGM thank CNPq-Brazil for partial financial support—Grants: 309063/2023-0 (RRC) and 307901/2022-0 (LGM). RRC and CAMM are grateful to FAPEMIG-Brazil (Grants APQ-00544- 23 and APQ-05218-23) for partial financial support. PJP acknowledges LAB-CCAM at ITA and CNPq (Grant 401565/2023-8).
@@ Appendix A Conserved current
where $\phi_{i}$ represents a colective set of fields with $i=1,2,\dots,N$.
@@
$\delta S=\int\left[\delta\left(d^{4}x\right)\mathcal{L}+d^{4}x\delta\mathcal{L}\right]$ (207)
@@
$+\frac{\partial\mathcal{L}}{\partial\left(\partial_{\nu}\partial_{\mu}\phi_{i}\right)}\bar{\delta}\left(\partial_{\nu}\partial_{\mu}\phi_{i}\right).$
@@
$+\partial_{\mu}{\left[\frac{\partial\mathcal{L}}{\partial\left(\partial_{\mu}\phi_{i}\right)}-2\partial_{\nu}\frac{\partial\mathcal{L}}{\partial\left(\partial_{\mu}\partial_{\nu}\phi_{i}\right)}\right]}\bar{\delta}\phi_{i},$
@@
By substituting Eqs. (A) and (209) in Eq. (208), we get
$\delta S$ $=$ $\int d^{4}x\left[{\frac{\delta\mathcal{L}}{\delta\phi_{i}}}\bar{\delta}\phi_{i}+\partial_{\mu}{\left[\frac{\partial\mathcal{L}}{\partial\left(\partial_{\mu}\phi_{i}\right)}-2\partial_{\nu}\frac{\partial\mathcal{L}}{\partial\left(\partial_{\mu}\partial_{\nu}\phi_{i}\right)}\right]}\bar{\delta}\phi_{i}\right.$ (213)
$\left.+\partial_{\mu}\left({\mathcal{L}\delta x^{\mu}}\right)+\partial_{\mu}\partial_{\nu}\left({\frac{\partial\mathcal{L}}{\partial\left(\partial_{\mu}\partial_{\nu}\phi_{i}\right)}}\bar{\delta}\phi_{i}\right)\right].$
Suppose that the total variation is generated by a transformation characterized by a set of parameters $\epsilon^{\alpha}$. In this case, the equations (201) and (202) assume the form
@@
$\bar{\delta}\phi_{i}\left(x\right)$ $=\epsilon^{a}\left[F_{i,a}\left(\phi,\partial\phi,\partial^{2}\phi\right)-A_{a}^{\sigma}\left(x\right)\partial_{\sigma}\phi_{i}\right].$
@@
Note that when the transformation considered is a symmetry of the action ($\delta S=0$), the solutions of the field equations $\phi_{i}^{sol}$ which satisfy the Euler-Lagrange equation
$\frac{\delta\mathcal{L}}{\delta\phi_{i}^{sol}}=0,$
lead to the conservation of the current:
$\left.\partial_{\mu}j_{a}^{\mu}\right|_{\phi_{i}=\phi_{i}^{sol}}=0.$ (215)

### v2

@@ IV Rotating ellipsoid with free precession
To analyze the impact of massive modes on rotational dynamics, we will focus our study specifically on the tensor mode within the oscillatory regime. In this case, by integrating Eqs. (194) and (195), we get the curves that are shown in Figs. 3 and 4. Figure 3 clearly shows that the frequency decreases with time, similar to what is predicted by GR. However, it also shows that the inclusion of a massive field moderates this frequency reduction, causing it to decrease more slowly as the mass $m_{\Psi}$ decreases. This happens because the reduction in $m_{\Psi}$ increases the destructive interference between the fields $h_{\mu\nu}$ and $\Psi_{\mu\nu}$, leading to a slower loss of energy by the system. A similar effect was observed in Alves et al. (2023) for point-like binary mass systems. The greater effectiveness of destructive interference with decreasing mass can be understood by recalling that $m_{\Psi}\propto\frac{1}{\alpha}$, where $\alpha$ is the coupling constant associated with the $C^{2}$ term in the action. Therefore, as $m_{\Psi}$ decreases, $\alpha$ increases, amplifying the impact of the massive spin-$2$ field.
@@ V Final comments
In this paper, we analyzed the emission of gravitational waves within the framework of quadratic gravity, including the contribution of longitudinal and transverse massive modes, extending the previous analysis presented in Ref. Alves et al. (2023). Firstly, we derived the wave equations for the massless ($\tilde{h}_{\alpha\beta}$) and massive ($\Phi,\Psi_{\alpha\beta}$) perturbations, followed by obtaining the associated conserved currents. Subsequently, we calculated the energy and angular momentum radiated through gravitational waves. Focusing on the quadrupole approximation, we examined two particular cases: a binary point-mass system in circular orbit, and a rotating ellipsoid with free precession.
@@ Acknowledgements
MFSA thanks FAPES (Brazil) for its support. The authors RRC and LGM thank CNPq-Brazil for partial financial support—Grants: 309063/2023-0 (RRC) and 307901/2022-0 (LGM). RRC and CAMM are grateful to FAPEMIG-Brazil (Grants APQ-00544- 23 and APQ-05218-23) for partial financial support. PJP acknowledges LAB-CCAM at ITA and CNPq (Grant 401565/2023-8). The authors are grateful to the referee whose comments helped to improve the paper.
@@ Appendix A Conserved current
where $\phi_{i}$ represents a collective set of fields with $i=1,2,\dots,N$.
@@
$\delta S$ $=\int\left[\delta\left(d^{4}x\right)\mathcal{L}+d^{4}x\delta\mathcal{L}\right]$
@@
$+\frac{\partial\mathcal{L}}{\partial\left(\partial_{\nu}\partial_{\mu}\phi_{i}\right)}\bar{\delta}\left(\partial_{\nu}\partial_{\mu}\phi_{i}\right).$ (204)
@@
$+\partial_{\mu}{\left[\frac{\partial\mathcal{L}}{\partial\left(\partial_{\mu}\phi_{i}\right)}-2\partial_{\nu}\frac{\partial\mathcal{L}}{\partial\left(\partial_{\mu}\partial_{\nu}\phi_{i}\right)}\right]}\bar{\delta}\phi_{i},$ (205)
@@
By substituting Eqs. (205) and (203) in Eq. (202), we get
$\delta S$ $=\int d^{4}x\left\{{\frac{\delta\mathcal{L}}{\delta\phi_{i}}}\bar{\delta}\phi_{i}+\partial_{\mu}{\left[\frac{\partial\mathcal{L}}{\partial\left(\partial_{\mu}\phi_{i}\right)}-2\partial_{\nu}\frac{\partial\mathcal{L}}{\partial\left(\partial_{\mu}\partial_{\nu}\phi_{i}\right)}\right]}\bar{\delta}\phi_{i}\right.$
$\left.+\partial_{\mu}\left({\mathcal{L}\delta x^{\mu}}\right)+\partial_{\mu}\partial_{\nu}\left({\frac{\partial\mathcal{L}}{\partial\left(\partial_{\mu}\partial_{\nu}\phi_{i}\right)}}\bar{\delta}\phi_{i}\right)\right\}.$ (207)
Suppose that the total variation is generated by a transformation characterized by a set of parameters $\epsilon^{a}$. In this case, the equations (196) and (197) assume the form
@@
$\Rightarrow\bar{\delta}\phi_{i}\left(x\right)=\epsilon^{a}\left[F_{i,a}\left(\phi,\partial\phi,\partial^{2}\phi\right)-A_{a}^{\sigma}\left(x\right)\partial_{\sigma}\phi_{i}\right].$
@@
Note that when the transformation considered is a symmetry of the action ($\delta S=0$), the solutions of the field equations $\phi_{i}^{\rm{sol}}$ which satisfy the Euler-Lagrange equation
$\frac{\delta\mathcal{L}}{\delta\phi_{i}^{\rm{sol}}}=0,$
lead to the conservation of the current, as follows:
$\left.\partial_{\mu}j_{a}^{\mu}\right|_{\phi_{i}=\phi_{i}^{\rm{sol}}}=0.$ (209)